STAFF RESOURCES

Minutes of July 9, 2003 meeting

Present:
I. Bogus, M. Brodman, A. DeChavez, S. Dematteo, M. Gadek, M. Ingegno, M. Katz, R. Krack, I. Loutchkina, J. Maguire, S. Peck, N. Reichardt, R. Smith, N. Sondhi, M. Spaventa, B. Stubbs, E. Suarez, N. Warren(recorder), D. Wilson
Absent:
R.Clarke, M. Herod, A. Martinez, E. McElroy, E. Moore, R. Riley, J. Teichmann, M. Tomaszewicz, P. Young

Agenda

ANNOUNCMENTS:

Functional group membership rosters have been finalized and now include staff from Collection Services and Access Services.

OLD BUSINESS:

CM MANUAL UPDATE- The manual is now complete, and copies were distributed to the group for review. Any changes can be reported to M. Brodman for updating, before final copies distributed. A web version of the manual is being developed for viewing on the NBL website.

STATISTICS REPORTING- Mary G. reminded the group to be timely in reporting monthly statistics. A discussion ensued on the overall importance of keeping accurate statistics, including the amount of time expended for shelving and shelf reading.

MISSING BOOK LIST- Lists will be distributed sometime in late July/early August. Results of the searches should be forwarded to the appropriate selectors.

SHIFTING UPDATES- Mary G. encouraged everyone to participate in the several shifting projects in progress across NBL, especially those members who have never shifted before. Interested members should email Mary with their availability.

-Kilmer is preparing to measure shelves in preparation for a shift in their stacks. Participation in this project was suggested for those new to shifting, to get an overview of the process. When weeding is complete in periodicals, shifting there will continue.

-LSM periodical shift is in progress. Measuring has been done and is being reviewed before the shift continues.

-Physics is planning a shift of their bound periodicals, in the coming weeks.

-After weeding has been completed at Music and Alexander reference will need shifts.

GOVERNMENT DOCUMENTS TRAINING- Because of the complexity of the classification system, participation was limited to select staff who the information would be most relevant to. Eddie, Rob and Brian recently completed training sessions, led by Nick. A need for a standardized format for the INT and UN sections, was discussed.

PRESERVATION DISPLAY- A committee, to be led by Mary B., was formed to begin work on a portable display and possible a Targum or RUTV advertisement. A discussion arouse about the standard food and drink policies across NBL.

NEW BUSINESS:

BINDERY TRAINING- One staff member from Alexander and one from LSM have been selected for training with Ian and Collection Services staff. They will spend two hours, twice a week for at least a month learning book repair techniques. Once the two are trained the can assist in the training of other staff.

INFORMATION SHARING:

Nicole/Chang- inventory of the collection continues. The Reference and Folio sections have been completed.

Roger/LSM- reports that Math continues the inventory of their collection, and Art will be weeding and shifting their Special Collections, some of which may be sent to Alex.

Stacy/Annex- materials being transferred to Annex from Douglas and Chemistry have stopped arriving. She reminded the group that the criteria for what can be sent to the Annex is on the NBL website, but should be sent to Collection Services first and not directly to her. Items moved during the renovation project are almost all replaced in their proper location.

Jamie/CS-LSM- due to shortages in staffing, she has been very busy.

Azon/CS-LSM- processing of bindery and serial check-in’s are up to date.

Neera/CS-Alex- has been busy converting 930 and 866 fields. Some Access Services staff will be trained for holdings conversions in the coming weeks.

Mary B./Alex- bound periodicals are arriving from Douglass, creating a need for more space. A shift is pending. It’s been unusually busy because of a shortage of voucher staff.

Shirley/ Alex- busy with recent increase of “In-Process” books.

Irina/Alex- has been helping Shirley with “In-Process” books, and is eager for new training opportunities.

Rob/Alex- shortage of voucher staff is keeping him busy with more routine tasks, cleaned up and organized storage space on 3rd floor and discovered a leak.

Eddie/LSM-continues work on the shift of LSM periodicals, a reorganization of staff shift scheduling may be necessary to make the project more time efficient, working with Collection Services to ensure the correct shelving of titles with acronyms.

Nick/LSM- a shift of the LSM microfiche collection is in progress and doing some weeding of material at Alex, to be sent to TSB for labeling.

Marty/LSM- busy processing RDS requests, while Drue is on vacation.

Ian/CS-NBL- backlog has been reduced with 735 books being sent to the bindery for repair, and a large number of brittle books have been boxed. Preparations are being made to train two Access Services staff members in some of the preservation techniques used by Collection Services.

Don/Kilmer- preparing for a shift in the stacks collection and has been busy with RDS while Asuncion is on vacation.

Matt/Douglass- a small shift, in the “H” section of the collection, is in progress.

Brian/Alex- creating a graphical interface for the Alexander collection. When a patron looks up a book in IRIS, a map will open showing them the exact location of the book in the stacks, including floor, column and row. It is designed for sections of the stacks and not individual books.

Megan/Physics- with the assistance of Howard Dess, weeding of the Physics permanent reserve collection continues and will hopefully be completed within the week.

Mary G./Douglass- construction project is moving along as scheduled, due to a problem with the exit gate at Douglass, the following information was discovered: When desensitizing books using the brick you should not slide the book back and forth, just slide the book one time in the direction of the arrow embossed on the brick, if there is no arrow just slide the book one time in one direction. By going back and forth you are desensitizing and the re-sensitizing, which results in the item not being completely desensitized and causes the gate alarm to sound.

NEXT MEETING: 8/13, location to be announced.



 
URL: http://www.libraries.rutgers.edu/rul/staff/groups/collection_mgt_group/minutes/collection_mgt_group_03_07_09.shtml
Website Feedback  |  Privacy Policy

© Copyright 1997-2013, Rutgers University Libraries   (Further Copyright Information)